OpenStudy (lena772):

If 4.35 grams of zinc metal react with 35.8 grams of silver nitrate, how many grams of silver metal can be formed and how many grams of the excess reactant will be left over when the reaction is complete? Show all of your work. unbalanced equation: Zn + AgNO3 “yields”/ Zn(NO3)2 + Ag

OpenStudy (lena772):

So far I have Zn + 2 AgNO3 = Zn(NO3)2 + 2 Ag 4.35 g / 65.39 g/mol=.0665 mol Zn .0665 * 2 = .133 mol AgNO3 required 35.8 g/ 169.87 g/mol=.2107 mol AgNO3 .2107 mol Ag .2107 * 107.868 g/mol=22.7277876 g Ag .2107/2 = .10535 mol Zn

OpenStudy (anonymous):

Still need help?

OpenStudy (lena772):

yes, what do i do now?

OpenStudy (anonymous):

Some of your steps were wrong.

OpenStudy (lena772):

ah, okay.

OpenStudy (anonymous):

Remember, the numbers in front of the compounds are ratios of MOLES. Not mass. You should have divided the masses (4.35 and 65.39) by their respective molar masses.

OpenStudy (lena772):

I did that in my steps.

OpenStudy (anonymous):

You divided 4.35 by 65.39. those are the masses of two different compounds.

OpenStudy (lena772):

65.39 is the mass of Zinc.

OpenStudy (lena772):

169.87 is the mass of silver nitrate.

OpenStudy (lena772):

So what did I do wrong?

OpenStudy (lena772):

I checked the source. I still don't see what you mean.

OpenStudy (lena772):

@halorazer

OpenStudy (anonymous):

Oh right. Duh. I'm an idiot. xD sorry. Let me try to work it out myself.

OpenStudy (lena772):

No, I make mistakes all the time. We're only human... Haha

OpenStudy (lena772):

is it 35.8-22.7277876=